# Zero-Shot Temporal Resolution Domain Adaptation for Spiking Neural Networks

### v2

@@ 4.4 A Simple Benchmark Method: Time-constant Adaptation Method
We now introduce a simple method that will serve as a benchmark for our proposed methods. This benchmark method is motivated by the explicit dependence of various neuron parameters on the sampling period in the popular LIF neuron and adaptive LIF neuron models, and has been utilized in the literature for the LIF neuron [he2020comparing]. Assume that the elements of the matrices $\bm{H}_{k,(S)}$ for $k=v,s,f,r$ belong to one of three groups:
@@
The time-constant adaptation methods obtains $\bm{H}_{k,(T)}$ for $k=v,s,f,r$, by replacing elements of $\bm{H}_{k,(S)}$ belonging to group i, ii, iii by:
@@
    rest of the terms are unchanged.

### v3

@@ 4.4 A Simple Benchmark Method: Time-constant Adaptation Method
We now introduce a simple method that will serve as a benchmark for our proposed methods. This benchmark method is motivated by the explicit dependence of various neuron parameters on the sampling period in the popular LIF neuron and adaptive LIF neuron models, and has been utilized in the literature for the LIF neuron [25]. Assume that the elements of the matrices $\bm{H}_{k,(S)}$ for $k=v,s,f,r$ belong to one of three groups:
@@ 2. $i$
The time-constant adaptation methods obtain $\bm{H}_{k,(T)}$ for $k=v,s,f,r$, by replacing elements of $\bm{H}_{k,(S)}$ belonging to the group i, ii, iii by:
@@
    rest of the terms are unchanged